# Model-Free Assessment of Simulator Fidelity via Quantile Curves

### v1

@@ 3 Theoretical Result
We introduce our method for constructing $\hat{V}$ and then provide theoretical guarantees, following up with an extension to directly compare simulators’ performance.
@@ 3.1 Methodology
Recall that our data $\mathcal{D}$ consists of $m$ scenarios; the $j$-th scenario has population parameters $p_{j}:=p(\psi_{j})$ and $q_{j}:=q(\psi_{j})$ in the real and simulator; $\hat{p}_{j}$ and $\hat{q}_{j}$ are their point estimates.
@@
    For each $j$, compute a confidence set $\mathcal{C}_{j}(\hat{p}_{j})$ on $p_{j}$, and calculate the pseudo-discrepancy $\hat{\Delta}_{j}:=\sup_{u\in\mathcal{C}_{j}\!\left(\hat{p}_{j}\right)}L\big(u,\hat{q}_{j}\big)$.
@@
    Denote $\hat{V}_{m}\!\left(\alpha\right):=$ the $\alpha$-quantile of $\{\hat{\Delta}_{j}\}_{j=1}^{m}$.
Intuitively, $\hat{\Delta}_{j}$ is a worst-case discrepancy for scenario $j$ over plausible values of $p_{j}$, and $\hat{V}_{m}(\alpha)$ is the empirical $\alpha$-quantile of these worst-case discrepancies. We will show in the next section that $\hat{V}_{m}\!\left(\alpha\right)$ is exactly the quantile function we are aiming for. Before that, we comment that the key design choice in our methodology is the construction of confidence set $\mathcal{C}_{j}\!\left(\cdot\right)$. Note that $\mathcal{C}_{j}\!\left(\cdot\right)$ is equivalent to the uncertainty set in robust optimization, which ensures that the true $p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j}\right)$ with a prescribed probability. The reason for introducing the uncertainty set will be discussed in the proof sketch. For now we illustrate how to construct these sets using off-the-shelf concentration bounds. In what follows we fix a coverage level $\gamma\in(0,1)$ (in Theorem 3.7 we take $\gamma=\tfrac{1}{2}$) and suppress $\gamma$ in the notation, writing $\mathcal{C}_{j}\!\left(\hat{p}_{j}\right):=\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)$. Consistent with our motivating example, we start with a multinomial setting.
@@
Suppose we have multinomial outcomes with $\Theta=\mathcal{P}^{d}:=\{u\in[0,1]^{d}:\sum_{i=1}^{d}u_{i}=1\}$ and denote KL divergence as $\mathrm{KL}(\cdot\|\cdot)$. Then
$\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j}\right))\geq\gamma,\text{ where }$ (3.1)
$\mathcal{C}_{j}\!\left(\hat{p}_{j}\right):=\Big\{$ $u\in\mathcal{P}^{d}:\mathrm{KL}(\hat{p}_{j}\|u)\leq\frac{d-1}{n_{j}}\log(\frac{2(d-1)}{\gamma})\Big\}.$
Bound (3.1) is a variant of Chernoff-Hoeffding Bound, see Lemma A.5 (Mardia et al., 2019). We also present examples for bounded outcomes, Bernoulli outcomes and general non-parametric distributions.
@@
Suppose outcome $Y$ is bounded between $[a,b]$ and $\hat{p}_{j}:=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}Y_{j,i}$ denote the sample mean for scenario $j$., then the set
$\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j}\right))\geq\gamma,\text{ where }$
$\mathcal{C}_{j}\!\left(\hat{p}_{j}\right):=\Big\{$ $u\in[a,b]:\ |u-\hat{p}_{j}|\leq(b-a)\sqrt{\tfrac{\log(2/\gamma)}{2n_{j}}}\Big\}.$
Example 3.3 (Bernoulli Confidence Set).
Suppose the outcome $Y$ is distributed according to a Bernoulli Distribution $Ber(p_{j})$. Then the set
$\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j}\right))\geq\gamma,\text{ where }$
$\mathcal{C}_{j}\!\left(\hat{p}_{j}\right):=\Big\{$ $u\in\mathbb{R}:\mathrm{KL}(\hat{p}_{j}\|u)\leq\frac{1}{n_{j}}\log(\frac{2}{\gamma})\Big\}.$
@@
For question $j$, let $\widehat{P}_{j}$ denote the empirical distribution from $n_{j}$ samples and assume that the true outcome $Y$ is $\sigma$-sub-Gaussian. Define the confidence set
$\mathcal{C}_{j}^{W_{1}}(\widehat{P}_{j}):=\Big\{Q:\ W_{1}(\widehat{P}_{j},Q)\leq r_{j}(n_{j},\gamma,\sigma)\Big\},$
where,
$r_{j}=\frac{512\sigma}{\sqrt{n_{j}}}+\sigma\sqrt{\frac{256e}{n_{j}}\log\frac{1}{1-\gamma}}.$
Then $\mathcal{C}_{j}^{W_{1}}(\widehat{P}_{j})$ covers the true distribution with probability at least $\gamma$. The above form is justified via concentration inequality proved in L.A. and Bhat 2022.
Next, we show that $\mathcal{C}_{j}$ can be efficiently computed for a given scenario $j$:
  • KL-Divergence: With the KL-based set $\mathcal{C}_{j}(\hat{p}_{j})=\{u\in\mathcal{P}^{d}:\ \mathrm{KL}(\hat{p}_{j}\|u)\leq r_{j}\}$ and loss $L(u,\hat{q}_{j})=\mathrm{KL}(u\|\hat{q}_{j})$, the pseudo-discrepancy is
    $\hat{\Delta}_{j}=\sup_{u\in\mathcal{C}_{j}(\hat{p}_{j})}\mathrm{KL}\big(u\|\hat{q}_{j}\big).$
    The maximizer for Bregman divergence lies on the boundary $\mathrm{KL}(\hat{p}_{j}\|u)=r_{j}$ and can be computed via a one-dimensional dual search for a Lagrange multiplier or other maximization techniques.
  • Wasserstein–1: For the $W_{1}$-based set $\mathcal{C}^{W_{1}}_{j}(\widehat{P}_{j})=\{u:W_{1}(u,\widehat{P}_{j})\leq r_{j}\}$ and loss $L(u,\widehat{Q}_{j})=W_{1}(u,\widehat{Q}_{j})$. Then, by the triangle inequality,
    $\hat{\Delta}_{j}=\sup_{u\in\mathcal{C}^{W_{1}}_{j}(\widehat{P}_{j})}W_{1}(u,\widehat{Q}_{j})\leq W_{1}(\widehat{P}_{j},\widehat{Q}_{j})+r_{j}.$
In both cases, $\hat{\Delta}_{j}$ reduces to a tractable, real-valued number per question, and other loss functions can be computed using similar techniques.
@@ 3.2 Calibrated Quantile Curve Theory
With the above methodology, we introduce two assumptions on which our theoretical guarantee relies and state the main theorem.
@@
Scenarios $\psi_{j}$ are drawn i.i.d. from $\Psi$. In addition, given $(\psi_{1},\dots,\psi_{m})$, the pairs $\mathcal{D}=\{(\mathcal{D}^{\mathrm{gt}}_{j},\mathcal{D}^{sim}_{j})\}_{j=1}^{m}$ are independent with $\mathcal{D}^{\mathrm{gt}}_{j}\perp\!\!\!\perp\mathcal{D}^{sim}_{j}$ conditional on $\psi_{j}$, and the new $(\psi,\mathcal{D}^{sim})$ is independent of $\mathcal{D}$.
@@
Suppose Assumption 3.1 and 3.2 hold. For any simulation sample size $k\in\mathbb{N}$, define the per-scenario discrepancy (unobservable) $\Delta_{j}^{(k)}$ and the pseudo-discrepancy (observable) $\hat{\Delta}_{j}^{(k)}$ by
$\Delta_{j}^{(k)}:=L\!\left(p_{j},\hat{q}_{j}\right),\qquad\hat{\Delta}_{j}^{(k)}:=\sup_{u\in\mathcal{C}_{j}\!\left(\hat{p}_{j}\right)}L\!\left(u,\hat{q}_{j}\right),$
where $\mathcal{C}_{j}\!\left(\hat{p}_{j}\right)\subset\Theta$ are data-driven compact confidence sets satisfying $\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j}\right)\big|\psi_{j},n_{j})\geq\tfrac{1}{2}.$ Then, for any $\alpha\in(0,1)$, with probability at least $1-\eta$ over $\mathcal{D}$, we have the following guarantee:
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}^{(k)}\leq\hat{V}_{m}\!\left(1-\tfrac{\alpha}{2}\right)\Big|\mathcal{D}\Big)\geq 1-\alpha-\frac{\varepsilon(\alpha,m,\eta)}{\sqrt{m}},$ (3.2)
where
$\varepsilon(\alpha,m,\eta)=$ $\sqrt{2\alpha\log\frac{2m}{\eta}+\frac{(\log\frac{2m}{\eta})^{2}+4\log\frac{2m}{\eta}}{m}}$
$+\frac{\log\tfrac{2m}{\eta}+2}{\sqrt{m}}+\sqrt{\frac{\log(4/\eta)}{2}}.$
In particular, the remainder is $O\big(\sqrt{(\log m)/m}\big)$, which vanishes as $m\to\infty$ for any $\alpha$.
We can rewrite the guarantee in Theorem 3.1 by the coverage level $\alpha$ rather than the tail parameter $1-\alpha$. In particular, for each fixed interior $\alpha$ and large $m$ this behaves as
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}^{(k)}\leq\hat{V}_{m}\Big(\tfrac{1+\alpha}{2}\Big)\Big|\mathcal{D}\Big)\approx\alpha-o_{m}(1).$ (3.3)
Heuristically, we can interpret the above formulation as follows: to get $\alpha$-level coverage for the true discrepancy distribution, we evaluate the empirical pseudo-quantile at the slightly more central index $(\alpha+1)/2$ (the finite-sample uncertainty), and we pay a vanishing $o_{m}(1)$ loss in coverage coming from having only $m$ scenarios (the scenario uncertainty). The complete proof is in Appendix B. We present the intuition of the proof below.
@@
For simplicity we ignore the dependence of $\Delta$ on $k$, as it will not affect the final theoretical guarantee. Suppose we have access to an oracle that returns the true $p_{j}$ for each scenario $j$, then this guarantees we have i.i.d. $\{\Delta_{j}\}^{m}_{j=1}$. Therefore, by adopting the well-known Dvoretzky-Kiefer-Wolfowitz (DKW) inequality on the $m$ scenarios, we can use the empirical quantile to upper-bound the true quantile function with probability $1-\eta$. However, we face two difficulties: (i) We do not have access to this oracle, (ii) each scenario has heterogeneous $n_{j}$ sample size.
To solve this problem, we wrap the uncertainty set $\mathcal{C}_{j}$ around the estimated parameter $\hat{p}_{j}$, which guarantees $\mathbb{P}(p_{j}\in\mathcal{C}_{j})\geq\tfrac{1}{2}$. Recall that we want to know the $\alpha$ quantile level of the actual quantile curve of $\Delta_{\psi}$ for any $\alpha$, and by taking $\sup$ over $\mathcal{C}_{j}$, we ensure that the pseudo-discrepancy exceeds the true discrepancy for that scenario with at least probability $\tfrac{1}{2}$, independently across scenarios. This allows us to infer the magnitude of the true discrepancy quantile from the pseudo-discrepancy quantile curve, since with probability $\geq\tfrac{1}{2}$ it provides an upper bound. To account for the part where the empirical quantile of pseudo-discrepancies doesn’t provide an upper bound to the true discrepancies, we shift to a more conservative tail index to align with the real sim-to-real tail. This accounts for one source of the correction terms, and along with the DKW correction terms as the other source, we obtain our theorem statement. ∎
@@
Note that we set $\gamma=\tfrac{1}{2}$ in Theorem 3.1, which can actually be changed to any $\gamma\in(0,1)$. The choice of $\gamma$ would directly affect the size of $\mathcal{C}_{j}$ and the tightness of the calibrated quantile curve, but will not affect the theoretical guarantee provided above.
[…]
We now turn to the benefits of obtaining a distributional level characterization of sim-to-real discrepancy. With this calibrated quantile, for an unseen scenario $\psi$ we can derive a confidence set for $p_{\psi}$ given a simulated $\hat{q}_{\psi}$. For a target confidence level $\bar{\alpha}\in(0,1)$, define

### v2

@@ 3 Theoretical Result
In this section, we construct $\hat{V}$, establish its theoretical guarantees, and conclude by discussing the benefits of distribution-level fidelity guarantees.
@@ 3.1 Methodology
Our dataset $\mathcal{D}$ consists of $m$ scenarios; the true parameter value (resp. finite sample estimate) for scenario $j$ is $p_{j}:=p(\psi_{j})$ (resp. $\hat{p}_{j}$); the simulator parameter value (resp. finite sample estimate) for scenario $j$ is $q_{j}:=q(\psi_{j})$ (resp. $\hat{q}_{j}$).
@@
    For each $j$, define a $\gamma_{j}$-confidence set $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$, i.e.
    $\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma_{j}\right))\geq\gamma_{j},$ (3.1)
    and the pseudo-discrepancy
    $\hat{\Delta}_{j}:=\sup_{u\in\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma_{j}\right)}L\big(u,\hat{q}_{j}\big).$ (3.2)
@@
    For $\alpha\in(0,1)$, let
    $\hat{V}_{m}\!\left(\alpha\right):=\text{the $\alpha$-quantile of}\{\hat{\Delta}_{j}\}_{j=1}^{m}.$ (3.3)
We will show that $\hat{V}_{m}\!\left(\alpha\right)$ is a calibrated quantile curve for the discrepancy distribution, up to explicit finite-sample corrections. Note that the key design choice in our method is the confidence set $\mathcal{C}_{j}\!\left(\cdot\right)$ that quantifies the uncertainty of $\hat{p}_{j}$. This uncertainty set is the key device that links per-scenario sampling error to distribution-level calibration. Next, we show how to construct these sets using off-the-shelf concentration bounds, starting with the multinomial setting.
@@
Suppose the outcomes are multinomial with parameter $p_{j}\in\Theta=\mathcal{P}^{d}:=\{u\in[0,1]^{d}:\sum_{i=1}^{d}u_{i}=1\}$, i.e., $Y_{j,i}\mid\psi_{j}\sim\mathrm{Categorical}(p_{j})$ and $\hat{p}_{j}$ is the empirical average vector from $n_{j}$ samples. Let $\mathrm{KL}(\cdot\|\cdot)$ denote the KL divergence. Then
$\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma_{j}\right):=\Big\{u\in\mathcal{P}^{d}:\mathrm{KL}(\hat{p}_{j}\|u)\leq\frac{d-1}{n_{j}}\ln(\frac{2(d-1)}{1-\gamma_{j}})\Big\}.$
is a $\gamma_{j}$-confidence set for $p_{j}$.
This result is a variant of Chernoff-Hoeffding Bound, see Lemma A.5 (Mardia et al., 2019).
@@
Suppose $Y_{j,i}\in[a,b]$ are i.i.d. with population mean $p_{j}:=\mathbb{E}[Y_{j,i}]$ for scenario $j$, and $\hat{p}_{j}:=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}Y_{j,i}$ is the sample mean. Then
$\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma_{j}\right):=\Big\{u\in[a,b]:|u-\hat{p}_{j}|\leq(b-a)\sqrt{\tfrac{\log(2/(1-\gamma_{j}))}{2n_{j}}}\Big\},$
is a $\gamma_{j}$-confidence set for $p_{j}$.
@@
Suppose the true outcome distribution for scenario $j$ is $P_{j}$ and samples are drawn as $Y_{j,1:n_{j}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}P_{j}$. Assume $P_{j}$ is $\sigma$-sub-Gaussian. Let $\widehat{P}_{j}$ denote the empirical distribution from $n_{j}$ samples. Then
$\mathcal{C}_{j}^{W_{1}}(\widehat{P}_{j},\gamma_{j}):=\Big\{Q:\ W_{1}(\widehat{P}_{j},Q)\leq r_{j}(n_{j},\gamma_{j},\sigma)\Big\},$
where
$r_{j}=\frac{512\sigma}{\sqrt{n_{j}}}+\sigma\sqrt{\frac{256e}{n_{j}}\log\frac{1}{1-\gamma_{j}}},$
is a $\gamma_{j}$-confidence set for the true distribution $P_{j}$.
The form of $\mathcal{C}_{j}$ is justified by the concentration inequality proved in L.A. and Bhat (2022).
Next, we show that $\hat{\Delta}_{j}$ can be efficiently computed for each scenario $j$.
  1. (a)
    KL-Divergence: For $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})=\{u\in\mathcal{P}^{d}:\ \mathrm{KL}(\hat{p}_{j}\|u)\leq r_{j}\}$ and loss $L(u,\hat{q}_{j})=\mathrm{KL}(u\|\hat{q}_{j})$, the pseudo-discrepancy
    $\hat{\Delta}_{j}=\sup_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})}\mathrm{KL}\big(u\|\hat{q}_{j}\big).$
    The maximizer lies on the boundary $\mathrm{KL}(\hat{p}_{j}\|u)=r_{j}$ and can be computed via standard maximization.
  2. (b)
    Wasserstein–1: For $\mathcal{C}^{W_{1}}_{j}(\widehat{P}_{j},\gamma_{j})=\{u:W_{1}(u,\widehat{P}_{j})\leq r_{j}\}$ and loss $L(u,\widehat{Q}_{j})=W_{1}(u,\widehat{Q}_{j})$, triangle inequality implies that
    $\hat{\Delta}_{j}=\sup_{u\in\mathcal{C}^{W_{1}}_{j}(\widehat{P}_{j},\gamma_{j})}W_{1}(u,\widehat{Q}_{j})\leq W_{1}(\widehat{P}_{j},\widehat{Q}_{j})+r_{j}.$
@@ 3.2 Calibrated Quantile Curve Theory
Our theoretical guarantee relies on the following two assumptions.
@@
The scenarios $\psi_{j}\sim\Psi$ are i.i.d. , $\mathcal{D}^{\mathrm{gt}}_{j}\perp\!\!\!\perp\mathcal{D}^{sim}_{j}$ conditioned on $\psi_{j}$, and the new $(\psi,\mathcal{D}^{sim})$ is independent of $\mathcal{D}$. In addition, the human sample sizes $\{n_{j}\}_{j=1}^{m}$ are treated as fixed (deterministic) design quantities.
@@
Before stating our main result, we consider the special case $\gamma_{j}\equiv\tfrac{1}{2}$. Then, in the asymptotic regime $m\to\infty$, our guarantee (3.6) below roughly implies the following: for any target quantile level $\tau\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}\leq\hat{V}_{m}\Big(\tfrac{1+\tau}{2}\Big)\Big|\mathcal{D}\Big)\geq\tau.$ (3.4)
In words, the $\tau$-quantile of the discrepancy $\Delta_{\psi}$ is approximated by the empirical pseudo-quantile curve at an inflated index $(1+\tau)/2$. This inflation induces an inherent conservativeness: for $\tau\approx 0$, the index $(1+\tau)/2$ is at least the median of the pseudo-discrepancy distribution, leading to substantial inflation in the lower tail.
In general, the $\gamma_{j}$’s can be different, and we obtain the following result:
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}\leq\hat{V}_{m}\big(\bar{\gamma}_{m}\tau+(1-\bar{\gamma}_{m})\big)\Big|\mathcal{D}\Big)\geq\tau,$ (3.5)
where $\bar{\gamma}_{m}=\tfrac{1}{m}\sum_{j=1}^{m}\gamma_{j}$. Now, to achieve $\tau$-level coverage we evaluate the empirical pseudo-quantile at an index that is affine in $\tau$ with slope $\bar{\gamma}_{m}$. One would ideally want $\bar{\gamma}_{m}=1$, or equivalently $\gamma_{j}\equiv 1$, to control the index inflation. However, increasing $\gamma_{j}$ typically enlarges the set $\mathcal{C}_{j}$, and thus, increases the values of the pseudo-discrepancies $\hat{\Delta}_{j}$ and shift the curve $\hat{V}_{m}$ upward. Thus, one needs to balance index adjustment with increased conservativeness of the pseudo-quantile curve itself.
This trade-off motivates choosing $\gamma_{j}$ adaptively as a function of the human sample size $n_{j}$. When $n_{j}$ is large, $p_{j}$ is typically more concentrated around $\hat{p}_{j}$; therefore, $\gamma_{j}$ can be set high without significantly enlarging $\mathcal{C}_{j}$. Therefore, we set $\gamma_{j}=g(n_{j})$, for an increasing function $g:\mathbb{R}_{+}\to[0,1]$. See Section 4 for numerical evidence illustrating the conservativeness of fixed $\gamma$ and the benefits of adaptive coverage. We are now ready to rigorously state the main theorem.
@@
Suppose Assumptions 3.13.2 hold and $\gamma_{j}\in(0,1)$ for each $j\in[m]$. Let $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$ denote any $\gamma_{j}$-confidence set (see (3.1)), $\hat{\Delta}_{j}$ denote the pseudo-discrepancy (see (3.2)), and let $\hat{V}_{m}$ denote the quantile function of the pseudo-discrepancy (see (3.3)). For $\alpha\in(0,1)$ let $\alpha]_{m}:=m^{-1}\lceil m\alpha\rceil\in\Big[\alpha,\alpha+\frac{1}{m}\Big]$ and set
$\alpha_{\mathrm{eff}}(\alpha):=\Big[\big(\bar{\gamma}_{m}-\varepsilon_{\mathrm{al}}(\alpha,\delta)\big)[\alpha]_{m}-b_{m}(\delta)\sqrt{[\alpha]_{m}}\Big]_{+},$
where $\bar{\gamma}_{m}:=\frac{1}{m}\sum_{j=1}^{m}\gamma_{j}$, $\varepsilon_{\mathrm{al}}(\alpha,\delta)=\sqrt{\frac{\log(3m/\delta)}{2m[\alpha]_{m}}}$, and $b_{m}(\delta)=\sqrt{\frac{1}{2m}\log\frac{3m}{\delta}}$. Then, with probability at least $1-\delta$ over $\mathcal{D}$, simultaneously for all $\alpha\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}^{(k)}\leq\hat{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha)\big)\Big|\mathcal{D}\Big)\geq 1-\alpha-\varepsilon_{m}(\delta),$ (3.6)
where $\varepsilon_{m}(\delta):=\sqrt{\frac{\log(6/\delta)}{2m}}+\tfrac{1}{m}$. In particular, the remainder terms $\varepsilon_{m},\varepsilon_{\mathrm{al}},$ and $b_{m}\to 0$ as $m\to\infty.$
Unlike standard conformal inference, which targets uncertainty quantification for future observable outcomes (e.g. responses), Theorem 3.1 provides uncertainty quantification for the latent, scenario-specific population parameter that governs outcome generation (e.g. conditional mean response). In other words, the guarantee operates at the level of underlying distributions, in contrast to single draws. We next provide a proof sketch.
@@
Suppose we have access to the true $p_{j}$ for each scenario $j$. Then $\{\Delta_{j}\}_{j=1}^{m}$ are i.i.d. , and the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality yields a uniform deviation bound for the empirical CDF of the discrepancy, which can be translated into upper bounds on its quantile function. In our setting, however, (i) $p_{j}$ is latent and (ii) each scenario is observed with a heterogeneous sample size $n_{j}$.
Our goal is to use empirical quantiles of the observable proxy $\{\hat{\Delta}_{j}\}$ to approximate the quantile function of the unobservable discrepancies $\{\Delta_{j}\}$. For simplicity, we fix $\gamma_{j}\equiv\gamma$. The key steps are as follows:
  • Let $t_{\alpha}$ denote the population $(1-\alpha)$-quantile of $\Delta_{j}$, and define the corresponding hardest scenarios $S_{\alpha}:=\{j:\Delta_{j}\geq t_{\alpha}\}$. By definition, $\tfrac{|S_{\alpha}|}{m}\approx\alpha$.
  • By construction, $\hat{\Delta}_{j}\geq\Delta_{j}$ holds with probability $\gamma$ for each $j$. Consequently, among indices in $S_{\alpha}$, at least a $\gamma$-fraction satisfy $\hat{\Delta}_{j}\geq t_{\alpha}$.
  • Hence, at least a $\gamma\alpha$-fraction of the pseudo-discrepancies exceed $t_{\alpha}$. In other words, the $(1-\alpha)$ quantile of $\Delta_{j}$ is dominated by the $(1-\gamma\alpha)$ quantile of $\hat{\Delta}_{j}$.
Therefore, for any quantile level $\tau:=1-\alpha$, taking the $(\gamma\tau+(1-\gamma))$-quantile of the pseudo-discrepancies yields an upper bound on the $\tau$-quantile of the oracle discrepancies.
This is why the quantile index must shift: to upper bound the oracle $\tau$-quantile of $\Delta$, we need to use a more conservative pseudo-quantile level so that the resulting pseudo-threshold is high enough to cover the target quantile despite the miscoverage. This index adjustment is the first source of conservativeness; the second source is the DKW deviation that relates empirical and population CDFs over finitely many scenarios. Combining these two ingredients yields the theorem. The complete proof is in Appendix B.1. ∎
@@
A subtle modeling choice is that Theorem 3.1 characterizes the distribution of $\Delta^{(k)}=L(p,\hat{q}^{(k)})$, where $\hat{q}^{(k)}$ is estimated using $k$ simulator samples, instead of the actual simulator parameter $q$. Conceptually, $\Delta^{(k)}$ is the correct target when one wants to characterize performance for a fixed query budget $k$. In contrast, $\Delta$ is more natural when simulator queries are cheap, and the goal also includes quantifying the simulator’s inherent bias. For brevity, we will, henceforth, ignore the $k$ superscript in following discussions. A proof to the latter case is presented in Corollary B.2, see Appendix B for details.
3.3 Benefits of Distribution-Level Characterization
We now turn to the benefits of obtaining a distributional-level characterization of sim-to-real discrepancy. For illustrative purposes we set $\bar{\gamma}_{m}=\tfrac{1}{2}$. With this calibrated quantile, for an unseen scenario $\psi$ we can derive a confidence set for $p_{\psi}$ given a simulated $\hat{q}_{\psi}$. For a target confidence level $\bar{\alpha}\in(0,1)$, define